# Exact diagonalization study of energy level statistics in harmonically confined interacting bosons

### v1

@@ III.3 Level number variance ${\Sigma^{2}(L)}$
The $\Delta_{3}(L)$ statistics and the level number variance $\Sigma^{2}(L)$ are related through the equation
[…]
From Eq. 6, we observe that $\Delta_{3}(L)$ is more uniform than $\Sigma^{2}(L)$, as it is the integrated version of the level number variance. This uniformity is evident in the plots of the $\Delta_{3}(L)$ statistics and the level number variance $\Sigma^{2}(L)$.
@@ III.4 Distribution of the ratio of consecutive level spacings
$P_{Poisson}(r)=\frac{2}{(1+r)^{2}}$ (8)
$P_{GOE}(r)=\frac{27}{4}\frac{r+r^{2}}{(1+r+r^{2})^{5/2}}.$ (9)
The advantage of the distribution of the ratios of consecutive level spacings $P(r)$ over the spacing distribution $P(s)$ is that we do not have to unfold the spectra. Thus, the distribution $P(r)$ is more useful quantity for comparison with experiments than the distribution of level spacings $P(s)$. The distribution $P(r)$ is used in many contexts of quantum many-body physics, such as in the investigation of many-body localization [10, 57, 58, 59], Bose-Hubbard models [26, 60] and the interaction driven metal-insulator crossover in the Hubbard model [61]. The average value of the gap ratio for the Poisson distribution is $\langle r\rangle_{Poisson}=2\ln(2)-1\approx 0.386$ [56], and for the GOE distribution, $\langle r\rangle_{GOE}\approx 0.530$ [56]. These averages of gap ratios can be obtained using the distributions given in Eqs. (8) and (9).
Table 1: Fitted values of the Brody parameter $b$ in the non-rotating case for different number of bosons in the moderate ($g_{2}=0.7301$) and the strong ($g_{2}=7.301$) interaction regimes.
Number of bosons, $N$ Brody parameter $b$
$g_{2}=0.7301$ $g_{2}=7.301$
12 0.33 0.53
16 0.16 0.85
20 0.04 0.51
Poisson 0
GOE 1
@@ IV Results & Discussions
Figure 1: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) in the moderate interaction regime with $g_{2}=0.7301$ for $N=12,16$ and 202020 bosons with total angular momentum $L_{z}=0$. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue dotted curve corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
We now present our numerical results on the spectral analysis employing the lowest 100100100 energy levels for the non-rotating case in Sec. IV.1 and for the rotating case in Sec. IV.2, considering both the moderate and the strong interaction regimes. To study the effect of the number of bosons $N$ on the statistical behavior of the energy levels, we vary the number of bosons as $N=12,16$, 202020. Our aim is to understand how the two-body interaction strength, rotation and variation in number of bosons affect the level statistics and hence the spectral correlations.
Table 2: The mean gap ratio $\langle r\rangle$ for different numbers of bosons in the non-rotating case in the moderate ($g_{2}=0.7301$) and the strong ($g_{2}=7.301$) interaction regimes.
Number of bosons, $N$ Mean gap ratio $\langle r\rangle$
$g_{2}=0.7301$ $g_{2}=7.301$
12 0.398 0.442
16 0.403 0.538
20 0.374 0.435
Poisson 0.386
GOE 0.530
@@ IV.1 Non-rotating case ($L_{z}=0$)
Figure 2: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) in the strong interaction regime with $g_{2}=7.301$ for $N=12,16$ and 202020 bosons with total angular momentum $L_{z}=0$. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue dotted curve corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
@@ IV.1.1 Short-range correlations
As the number of bosons increases from $N=12,16$ to 202020, the NNSD aligns more towards Poisson distribution as shown in Figs. 1(a), 1(b), and 1(c), respectively (the upper panel). A more quantitative measure of the behavior is given by the Brody parameter $b$ obtained by fitting Eq. (2). In Table 1, we present the Brody parameter $b$ for $N=12,16$ and 202020, which in the moderate interaction regime (with $g_{2}=0.7301$) is found to decrease systematically with increase in number of bosons (third subcolumn). For $N=12$, the Brody parameter is $b=0.33$, signifying small correlations between the energy levels. However, as the number of bosons increases to $N=16$, the Brody parameter decreases to $b=0.16$ and for $N=20$, the Brody parameter further decreases to $b=0.04$, indicating a Poissonian distribution where the level repulsion is minimum. The observed behavior implies that an increase in number of bosons $N$ leads to decrease in spectral correlations. This shows that, in this regime, the energy levels remain largely uncorrelated, a characteristic feature of a regular system.
The results for the consecutive ratios of level spacings distribution $P(r)$ with increasing number of bosons $N=12$, 161616 and 202020 are presented in Figs. 1(d), 1(e) and 1(f), respectively (the lower panel). The corresponding values of the mean gap ratio $\langle r\rangle$ are summarized in the third subcolumn of Table 2. For $N=12$, the calculated mean gap ratio is $\langle r\rangle=0.398$. As the number of bosons increases to $N=16$, the mean gap ratio shows a marginal increase to $\langle r\rangle=0.403$. Upon further increase in the number of bosons to $N=20$, the mean gap ratio becomes $\langle r\rangle=0.374$. These values of $\langle r\rangle$ are in close agreement with the theoretical value of $\langle r\rangle\approx 0.386$ for a Poissonian distribution, indicating that the system primarily features regular behavior with no significant spectral correlations.
Table 3: Fitted values of the Brody parameter $b$ in the rotating single-vortex state $L_{z}=N$ for different number of bosons in the moderate ($g_{2}=0.7301$) and the strong ($g_{2}=7.301$) interaction regimes.
Number of bosons, $N$ $L_{z}=N$ Brody parameter $b$
$g_{2}=0.7301$ $g_{2}=7.301$
12 12 0.52 1.18
16 16 0.64 0.98
20 20 0.64 1.07
@@
Strong interaction regime
The results on NNSD $P(s)$ for varying numbers of bosons $N=12,16,20$ are presented in Figs. 2(a)2(c), in the strong interaction regime ($g_{2}=7.301$). For $N=12$, the NNSD follows the GOE distribution with Brody parameter $b=0.53$, indicating small spectral correlations leading to a weakly chaotic behavior, as shown in Fig. 2(a). This suggests that while signatures of chaotic behavior are present, a degree of regularity persists within the system. As the number of bosons increases to $N=16$, the spectral correlation between energy levels enhances, as evidenced by increase in Brody parameter to the value $b=0.85$, shown in Fig. 2(b). This considerable increase in Brody parameter implies that the system has developed a significantly chaotic behavior, exhibiting strong level repulsion, a characteristic of GOE statistics. However, for $N=20$, the Brody parameter decreases to $b=0.51$, as seen in Fig. 2(c), indicating that the system returns to a weakly chaotic regime with revival of some degree of regularity within the system. The corresponding values of the Brody parameter $b$ are summarized in the fourth subcolumn of Table 1.
The distribution of the ratio of consecutive level spacings $P(r)$ for the increasing number of bosons $N=12$, 161616 to 202020 is shown in Figs. 2(d)2(f). For $N=12$, the $P(r)$ distribution begins to deviate from Poissonian statistics and shows characteristics similar to those of GOE, as seen in Fig. 2(d). The corresponding mean gap ratio is $\langle r\rangle=0.442$, indicating an intermediate regime in which the system exhibits signatures of weak chaotic behavior with some degree of regularity. As the number of bosons increases to $N=16$, the $P(r)$ distribution aligns closely with GOE distribution, as seen in Fig. 2(e). The calculated mean gap ratio in this case is $\langle r\rangle=0.538$, which is close to the theoretical value of $\langle r\rangle\approx 0.530$ for GOE distribution. This indicates the emergence of strong spectral correlations, suggesting significant quantum-chaotic behavior in the system. As the number of bosons increases further to $N=20$, the $P(r)$ distribution returns to align with GOE distribution with revival of some degree of regularity in the system, as shown in Fig. 2(f). The calculated mean gap ratio decreases to $\langle r\rangle=0.435$, implying a weakening of chaotic behavior. The values of mean gap ratio $\langle r\rangle$ corresponding to each number of bosons $N$ are summarized in the fourth subcolumn of Table 2.
Table 4: The mean gap ratio $\langle r\rangle$ for different numbers of bosons in the rotating single-vortex state $L_{z}=N$ in the moderate ($g_{2}=0.7301$) and the strong ($g_{2}=7.301$) interaction regimes.
Number of bosons, $N$ $L_{z}=N$ Mean gap ratio $\langle r\rangle$
$g_{2}=0.7301$ $g_{2}=7.301$
12 12 0.498 0.555
16 16 0.494 0.505
20 20 0.496 0.549
@@
(a) $N=12,16$ and 202020, $g_{2}=0.7301$
(b) $N=12,16$ and 202020, $g_{2}=0.7301$
(c) $N=12,16$ and 202020, $g_{2}=7.301$
(d) $N=12,16$ and 202020, $g_{2}=7.301$
Figure 3: (Color online) The spectral average $\langle\Delta_{3}(L)\rangle$ vs $L$ and the level number variance ${\Sigma^{2}(L)}$ vs $L$ are presented for moderate interaction regime with $g_{2}=0.7301$ (upper panel) and strong interaction regime with $g_{2}=7.301$ (lower panel) for different number of bosons $N=12,16$ and 202020 with total angular momentum $L_{z}=0$. The blue circle, green square and red diamond lines are our numerical results of $\langle\Delta_{3}(L)\rangle$ (${\Sigma^{2}(L)}$) for $N=12,16$ and 202020, respectively, for the lowest 100 energy levels . For reference, we have also drawn $\langle\Delta_{3}(L)\rangle$ (${\Sigma^{2}(L)}$), as a function of $L$, corressponding to the Poisson distribution (black dashed line) and the GOE distribution (magenta dash-dot line).
@@ IV.1.2 Long-range correlations
The spectral average $\Delta_{3}(L)$ for moderate interaction regime is calculated over the interval $L=2$ to 606060, where $L$ denotes the length of the energy interval. We linearly fit our numerical data for $\Delta_{3}(L)$ to obtain the slope of $\Delta_{3}(L)$ vs $L$, upto $L=10$ and compare it with the theoretical value $\Delta_{3}(L)=\frac{L}{15}=0.0667L$ of Poisson distribution.
For $N=12$, the $\Delta_{3}(L)$ statistic exhibits a linear increase with a slope of 0.0467 up to $L=10$, beyond which it saturates to a constant value, as shown in Fig. 3(a). A similar trend is observed for $N=16$ and 202020, where the corresponding slopes are 0.0569 and 0.0634, respectively, as shown in Fig. 3(a). Thus, as the number of bosons increases, the behavior of $\Delta_{3}(L)$ agrees more with Poisson distribution upto energy interval $L=10$, indicating decrease in spectral correlations in the system. At large values of $L$, the saturation of $\Delta_{3}(L)$, as seen in Fig. 3(a), is consistent with the argument given by Berry [64].
Another indicator of long-range spectral correlation is the level number variance $\Sigma^{2}(L)$. We perform a linear fit to our numerical data to determine the slope of $\Sigma^{2}(L)$ vs $L$, up to $L=2$ and compare it with the theoretical value $\Sigma^{2}(L)=L$ of Poisson distribution. For $N=12,16$ and 202020, the corresponding slope values are 0.6615,0.75610.6615,0.75610.6615 , 0.7561 and 0.94020.94020.9402, respectively, indicating an increasing alignment with Poisson distribution, which has slope of 111, as shown in Fig. 3(b). The increasing slope with increasing number of bosons $N$ suggests a reduction in spectral correlations in the system. Thus, $\Sigma^{2}(L)$, on the average, closely follows Poisson distribution, up to $L=2$, while for $L>2$, it exhibits noticeable deviation, as shown in Fig. 3(b).
@@
The spectral average $\Delta_{3}(L)$ for different number of bosons $N$, in strong interaction regime, is shown in Fig. 3(c). Our numerical results demonstrates that the $\Delta_{3}(L)$ statistic follows the GOE distribution up to a certain range of the energy interval $L$, beyond which it saturates to a constant value $\Delta_{\infty}=0.543$ (for $N=20$). Notably, this saturation occurs at a significantly lower value compared to the moderate interaction case, where $\Delta_{\infty}=1.48$ for the same number of bosons [64]. As the number of bosons increases from $N=12,16$, 202020, the saturation value of $\Delta_{3}(L)$ at large $L$ also increases, as shown in Fig. 3(c).
@@
IV.2 Rotating case
We focus on the rotating single-vortex state $L_{z}=N$ for number of bosons $N=12,16$ and 202020 for both the short-range and the long-range correlations.
@@ IV.2 Rotating case
Figure 4: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) for moderate interaction regime with $g_{2}=0.7301$ in the single-vortex state $L_{z}=N$ where $N=12,16$ and 202020. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue line corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution, and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
IV.2.1 Short-range correlations
The distributions $P(s)$ and $P(r)$ for various values of $L_{z}=N$ corresponding to number of bosons $N=12,16$ and 202020, are presented in Fig. 4 for the moderate interaction regime and in Fig. 6 for the strong interaction regime.
Moderate interaction regime
For number of bosons $N=12,16$, 202020 and for the single vortex state $L_{z}=N$, the NNSD $P(s)$ distribution exhibits GOE behavior with Brody parameter $b=0.52,0.64$ and 0.640.640.64, respectively, as shown in Figs. 4(a)-4(c). This indicates that these systems exist in a weakly chaotic regime, retaining substantial degree of regularity.
The gap ratio distribution $P(r)$ for $N=12,16$ and 202020 in the single vortex state $L_{z}=N$ are presented in Figs. 4(d)-4(f) with mean gap ratio $\langle r\rangle=0.498,0.494$ and 0.4960.4960.496, respectively, indicating that these systems remain in a weakly chaotic regime.
Strong interaction regime
The gap ratio distribution $P(r)$ for $N=12,16,20$ in the single vortex state $L_{z}=N$ follows the GOE distribution with mean gap ratio $\langle r\rangle=0.555,0.505$ and 0.5490.5490.549, respectively, as shown in Figs. 5(d)-5(f). This indicate that these systems show strong signature of quantum chaos.
The Brody parameter $b$ and the mean gap ratio $\langle r\rangle$ values for each of the number of bosons $N$ in the single vortex state $L_{z}=N$ with moderate and strong interaction regimes are listed in Table 3 and 4.
IV.2.2 Long-range correlations
Moderate interaction regime
The spectral average $\Delta_{3}(L)$ for $N=12,16,20$ in the single vortex state $L_{z}=N$ is shown in Fig. 6(a). It is observed that for small energy intervals $L$, the $\Delta_{3}(L)$ statistic exhibits GOE like behavior, while for large $L$, it approaches a saturation value. The saturated values for $N=12,16$ and 202020 are found to be $\Delta_{\infty}$= 0.9160.9160.916, 0.4030.4030.403 and 0.6570.6570.657, respectively. The value of $\Delta_{\infty}$ is modulated by the number of bosons $N$ in the moderate interaction regime which may be attributed to the interplay between the number of bosons and the interaction strength.

### v2

@@ III.3 Level number variance $\Sigma^{2}(L)$}
The $\Delta_{3}(L)$ statistic and the level number variance $\Sigma^{2}(L)$ are related through the equation
$\Delta_{3}(L)=\frac{2}{L^{4}}\int_{0}^{L}(L^{3}-2L^{2}\epsilon+\epsilon^{3})\Sigma^{2}(\epsilon)d\epsilon.$ (8)
From Eq. (8), we observe that $\Delta_{3}(L)$ is more uniform than $\Sigma^{2}(L)$, as it is the integrated version of the level number variance. This uniformity is evident in the plots of the $\Delta_{3}(L)$ statistic and the level number variance $\Sigma^{2}(L)$.
@@ III.4 Distribution of the ratio of consecutive level spacings
$P_{\mathrm{Poisson}}(r)=\frac{2}{(1+r)^{2}}$ (10)
$P_{\mathrm{GOE}}(r)=\frac{27}{4}\frac{r+r^{2}}{(1+r+r^{2})^{5/2}}.$ (11)
The advantage of the distribution of the ratios of consecutive level spacings $P(r)$ over the spacing distribution $P(s)$ is that we do not have to unfold the spectra. Thus, the distribution $P(r)$ is more useful quantity for comparison with experiments than the distribution of level spacings $P(s)$. The distribution $P(r)$ is used in many contexts of quantum many-body physics, such as in the investigation of many-body localization [11, 59, 60, 61], Bose-Hubbard models [27, 62] and the interaction driven metal-insulator crossover in the Hubbard model [63]. The average value of the gap ratio for the Poisson distribution is $\langle r\rangle_{\mathrm{Poisson}}=2\ln(2)-1\approx 0.386$ [58], and for the GOE distribution, $\langle r\rangle_{\mathrm{GOE}}\approx 0.530$ [58]. These averages of gap ratios can be obtained using the distributions given in Eqs. (10) and (11).
@@ IV Results & Discussions
Figure 1: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) in the moderate interaction regime with $g_{2}=0.3669$ for $N=12,16$ and $20$ bosons with total angular momentum $L_{z}=0$. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue dotted curve corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
We now present our numerical results on the spectral analysis employing the lowest $100$ energy levels for the non-rotating case in Sec. IV.1 and for the rotating case in Sec. IV.2, considering both the moderate and the strong interaction regimes. To study the effect of the number of bosons $N$ on the statistical behavior of the energy levels, we vary the number of bosons as $N=12,16$, 2020. Our aim is to understand how the variation in two-body interaction, rotation and number of bosons affect the level statistics and hence the spectral correlations.
Table 1: Fitted values of the Brody parameter $b$ in the non-rotating case for different number of bosons in the moderate ($g_{2}=0.3669$) and the strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Brody parameter $b$
$g_{2}=0.3669$ $g_{2}=3.669$
12 0.33 0.53
16 0.16 0.85
20 0.04 0.51
Poisson 0
GOE 1
@@ IV.1 Non-rotating case ($L_{z}=0$)
Figure 2: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) in the strong interaction regime with $g_{2}=3.669$ for $N=12,16$ and $20$ bosons with total angular momentum $L_{z}=0$. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue dotted curve corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
(a) $N=12,16$ and $20$, $g_{2}=0.3669$
(b) $N=12,16$ and $20$, $g_{2}=0.3669$
(c) $N=12,16$ and $20$, $g_{2}=3.669$
(d) $N=12,16$ and $20$, $g_{2}=3.669$
Figure 3: (Color online) The spectral average $\langle\Delta_{3}(L)\rangle$ vs $L$ and the level number variance $\Sigma^{2}(L)$} vs $L$ are presented for moderate interaction regime with $g_{2}=0.3669$ (upper panel) and strong interaction regime with $g_{2}=3.669$ (lower panel) for different number of bosons $N=12,16$ and 2020 with total angular momentum $L_{z}=0$. The blue circle, green square and red diamond lines are our numerical results of $\langle\Delta_{3}(L)\rangle$ ($\Sigma^{2}(L)$}) for $N=12,16$ and $20$, respectively, for the lowest 100 energy levels . For reference, we have also drawn $\langle\Delta_{3}(L)\rangle$ ($\Sigma^{2}(L)$}), as a function of $L$, corressponding to the Poisson distribution (black dashed line) and the GOE distribution (magenta dash-dot line).
@@ IV.1.1 Short-range correlations
As the number of bosons increases from $N=12,16$ to $20$, the NNSD aligns more towards Poisson distribution as shown in Figs. 1(a), 1(b), and 1(c), respectively (the upper panel). A more quantitative measure of the behavior is given by the Brody parameter $b$ obtained by fitting Eq. (4). In Table 1, we present the Brody parameter $b$ for $N=12,16$ and $20$, which in the moderate interaction regime (with $g_{2}=0.3669$) is found to decrease systematically with increase in number of bosons (third subcolumn). For $N=12$, the Brody parameter is $b=0.33$, signifying small correlations between the energy levels. However, as the number of bosons increases to $N=16$, the Brody parameter decreases to $b=0.16$ and for $N=20$, the Brody parameter further decreases to $b=0.04$, indicating a Poissonian distribution where the level repulsion is minimum. The observed behavior implies that an increase in number of bosons $N$ leads to decrease in spectral correlations. This shows that, in this regime, the energy levels remain largely uncorrelated, a characteristic feature of a regular system.
Table 2: The mean gap ratio $\langle r\rangle$ with estimated error for different number of bosons in the non-rotating case in the moderate ($g_{2}=0.3669$) and the strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Mean gap ratio $\langle r\rangle$
$g_{2}=0.3669$ $g_{2}=3.669$
12 0.398 $\pm$ 0.027 0.442 $\pm$ 0.028
16 0.403 $\pm$ 0.030 0.538 $\pm$ 0.028
20 0.374 $\pm$ 0.028 0.435 $\pm$ 0.027
Poisson 0.386
GOE 0.530
Table 3: Fitted values of the Brody parameter $b$ in the rotating single-vortex state $L_{z}=N$ for different number of bosons in the moderate ($g_{2}=0.3669$) and the strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Brody parameter $b$
$g_{2}=0.3669$ $g_{2}=3.669$
12 0.52 1.18
16 0.64 0.98
20 0.64 1.07
@@
The results for the consecutive ratios of level spacings distribution $P(r)$ with increasing number of bosons $N=12$, $16$ and $20$ are presented in Figs. 1(d), 1(e) and 1(f), respectively (the lower panel). The corresponding values of the mean gap ratio $\langle r\rangle$ are summarized in the third subcolumn of Table 2. For $N=12$, the calculated mean gap ratio is $\langle r\rangle=0.398$. As the number of bosons increases to $N=16$, the mean gap ratio shows a marginal increase to $\langle r\rangle=0.403$. Upon further increase in the number of bosons to $N=20$, the mean gap ratio becomes $\langle r\rangle=0.374$. These values of $\langle r\rangle$ are in close agreement with the theoretical value of $\langle r\rangle\approx 0.386$ for a Poissonian distribution, indicating that the system primarily features regular behavior with no significant spectral correlations. The mean gap ratio for the two subsets of lowest energy levels—the lower 1-50 and the higher 51–100 levels—is shown in fourth column of Table 5. It is observed that the lower levels (115050) exhibit regular behavior, as the mean gap ratio is close to the theoretical Poisson value of $0.386$. The higher levels (5151100100) also show regular behavior with a slightly larger mean gap ratio compared to the lower levels.
Table 4: The mean gap ratio $\langle r\rangle$ with estimated error for different number of bosons in the rotating single-vortex state $L_{z}=N$ in the moderate ($g_{2}=0.3669$) and the strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Mean gap ratio $\langle r\rangle$
$g_{2}=0.3669$ $g_{2}=3.669$
12 0.498 $\pm$ 0.027 0.555 $\pm$ 0.026
16 0.494 $\pm$ 0.028 0.505 $\pm$ 0.027
20 0.496 $\pm$ 0.025 0.549 $\pm$ 0.023
@@
Strong interaction regime
The results on NNSD $P(s)$ for varying numbers of bosons $N=12,16,20$ are presented in Figs. 2(a)2(c), in the strong interaction regime ($g_{2}=3.669$). For $N=12$, the NNSD follows the GOE distribution with Brody parameter $b=0.53$, indicating small spectral correlations leading to a weakly chaotic behavior, as shown in Fig. 2(a). This suggests that while signatures of chaotic behavior are present, a degree of regularity persists within the system. As the number of bosons increases to $N=16$, the spectral correlation between energy levels enhances, as evidenced by increase in Brody parameter to the value $b=0.85$, shown in Fig. 2(b). This considerable increase in Brody parameter implies that the system has developed a significantly chaotic behavior, exhibiting strong level repulsion, a characteristic of GOE statistics. However, for $N=20$, the Brody parameter decreases to $b=0.51$, as seen in Fig. 2(c), indicating that the system returns to a weakly chaotic regime with revival of some degree of regularity within the system. The corresponding values of the Brody parameter $b$ are summarized in the fourth subcolumn of Table 1.
Table 5: The mean gap ratio with estimated error for the two subsets of energy levels—the lower 1–50 and the higher 51–100 levels—is calculated for different number of bosons in the non-rotating case ($L_{z}=0$) in both moderate ($g_{2}=0.3669$) and strong ($g_{2}=3.669$) interaction regimes.
Number of bosons, NN Number of levels Mean gap ratio $\langle r\rangle$
$g_{2}=0.3669$ $g_{2}=3.669$
12 1–50 0.363 $\pm$ 0.040 0.411 $\pm$ 0.042
51–100 0.432 $\pm$ 0.037 0.472 $\pm$ 0.038
16 1–50 0.331 $\pm$ 0.039 0.572 $\pm$ 0.043
51–100 0.478 $\pm$ 0.044 0.499 $\pm$ 0.035
20 1–50 0.338 $\pm$ 0.038 0.492 $\pm$ 0.041
51–100 0.423 $\pm$ 0.042 0.382 $\pm$ 0.035
Poisson 0.386
GOE 0.530
The distribution of the ratio of consecutive level spacings $P(r)$ for the increasing number of bosons $N=12$, $16$ to $20$ is shown in Figs. 2(d)2(f). For $N=12$, the $P(r)$ distribution begins to deviate from Poissonian statistics and shows characteristics similar to those of GOE, as seen in Fig. 2(d). The corresponding mean gap ratio is $\langle r\rangle=0.442$, indicating an intermediate regime in which the system exhibits signatures of weak chaotic behavior with some degree of regularity. As the number of bosons increases to $N=16$, the $P(r)$ distribution aligns closely with GOE distribution, as seen in Fig. 2(e). The calculated mean gap ratio in this case is $\langle r\rangle=0.538$, which is close to the theoretical value of $\langle r\rangle\approx 0.530$ for GOE distribution. This indicates the emergence of strong spectral correlations, suggesting significant quantum-chaotic behavior in the system. As the number of bosons increases further to $N=20$, the $P(r)$ distribution returns to align with GOE distribution with revival of some degree of regularity in the system, as shown in Fig. 2(f). The calculated mean gap ratio decreases to $\langle r\rangle=0.435$, implying a weakening of chaotic behavior. The values of mean gap ratio $\langle r\rangle$ corresponding to each number of bosons $N$ are summarized in the fourth subcolumn of Table 2. The mean gap ratio for the lower (115050) and higher (5151100100) subsets of energy levels is summarized in the fifth column of Table 5. The lower (115050) part of the spectrum exhibits chaotic behavior consistent with GOE distribution (mean gap ratio, $\langle r\rangle\approx 0.530$). The higher (5151100100) levels retain signatures of chaoticity, except for $N=20$, where the mean gap ratio shows deviation from GOE distribution to align with Poisson distribution.
@@ IV.1.2 Long-range correlations
The spectral average $\Delta_{3}(L)$ for moderate interaction regime is calculated over the interval $L=2$ to $60$, where $L$ denotes the length of the energy interval. We linearly fit our numerical data for $\Delta_{3}(L)$ to obtain the slope of $\Delta_{3}(L)$ vs $L$, upto $L=10$ and compare it with the theoretical value $\Delta_{3}(L)=\frac{L}{15}=0.0667L$ of Poisson distribution.
For $N=12$, the $\Delta_{3}(L)$ statistic exhibits a linear increase with a slope of 0.0467 up to $L=10$, beyond which it saturates to a constant value, as shown in Fig. 3(a). A similar trend is observed for $N=16$ and $20$, where the corresponding slopes are 0.0569 and 0.0634, respectively, as shown in Fig. 3(a). Thus, as the number of bosons increases, the behavior of $\Delta_{3}(L)$ agrees more with Poisson distribution upto energy interval $L=10$, indicating decrease in spectral correlations in the system. At large values of $L$, the $\Delta_{3}(L)$ statistic saturates to $\Delta_{\infty}=1.74$, $1.50$, and $1.48$ for $N=12$, $16$, and $20$, respectively, as shown in Fig. 3(a). This behavior is consistent with Berry’s argument [66] that $\Delta_{3}(L)$ approaches a constant value in the large $L$ limit.
Another indicator of long-range spectral correlation is the level number variance $\Sigma^{2}(L)$. We perform a linear fit to our numerical data to determine the slope of $\Sigma^{2}(L)$ vs $L$, up to $L=2$ and compare it with the theoretical value $\Sigma^{2}(L)=L$ of Poisson distribution. For $N=12,16$ and $20$, the corresponding slope values are $0.6615,0.7561$ and $0.9402$, respectively, indicating an increasing alignment with Poisson distribution, which has slope of $1$, as shown in Fig. 3(b). The increasing slope with increasing number of bosons $N$ suggests a reduction in spectral correlations in the system. Thus, $\Sigma^{2}(L)$, on the average, closely follows Poisson distribution, up to $L=2$, while for $L>2$, it exhibits noticeable deviation, as shown in Fig. 3(b).
@@
The spectral average $\Delta_{3}(L)$ for different number of bosons $N$, in strong interaction regime, is shown in Fig. 3(c). Our numerical result demonstrates that the $\Delta_{3}(L)$ statistic follows the GOE distribution up to a certain range of the energy interval $L$, beyond which it saturates to a constant value $\Delta_{\infty}=0.321,0.475$ and $0.543$ for $N=12,16$ and $20$, respectively. Notably, this saturation for the strong interaction occurs at a significantly lower value compared to the moderate interaction case, where $\Delta_{\infty}=1.74,1.50$ and $1.48$ for the respective number of bosons. This is consistent with Berry’s argument that integrable systems exhibit higher saturated $\Delta_{3}$ values than non-integrable (chaotic) ones [66]. For example, in the non-rotating case with $N=12$, the saturated values of $\Delta_{3}$ are found to be 1.74 and 0.321 for the moderate and strong interaction regimes, respectively. This suggests that the system is integrable in the moderate interaction regime and becomes non-integrable in the strong interaction regime. Similar trend is observed for $N=16$ and $N=20$ across both interaction regimes. As the number of bosons increases from $N=12,16$, 2020, the saturation value of $\Delta_{3}(L)$ at large $L$ also increases, as shown in Fig. 3(c).
@@
Figure 4: (Color online) The nearest-neighbor spacing distribution $P(s)$ (upper panel) and the distribution of the ratio of consecutive level spacings $P(r)$ (lower panel) for moderate interaction regime with $g_{2}=0.3669$ in the single-vortex state $L_{z}=N$ where $N=12,16$ and $20$. The histogram in each graphs represents our numerical result for the lowest 100 energy levels. The blue line corresponds to the Poisson distribution, the orange dashed curve to the GOE distribution, and the green dash-dotted curve to the Brody distribution with fitting parameter $b$.
IV.2 Rotating case
We focus on the rotating single-vortex state $L_{z}=N$ for number of bosons $N=12,16$ and $20$ for both the short-range and the long-range correlations.